OpenStudy (anonymous):

Indicate which property is illustrated in Step 6. Step 1 8 - 6x - 8 + 7x = (8 - 6x) + (-8 + 7x) Step 2 = (8 - 6x) + (7x - 8) Step 3 = 8 + (-6x + 7x) - 8 Step 4 = 8 + (-6 + 7)x - 8 Step 5 = 8 + x - 8 Step 6 = x + 8 - 8 Step 7 = x + 0 Step 8 = x distributive identity commutative associative

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

I believe this is identity property but I'm not sure

OpenStudy (triciaal):

I think commutative only rearranging the order a + b + c = b + a + c 8 + x - 8 Step 6 = x + 8 - 8